3. Consider a reaction in which hydrogen (H2) and oxygen (02) react to form water (H2) By mass, the mixture of reactants is 1 part hydrogen and 6 parts oxygen Write the chemical reaction showing reactant and products including any left over oxygen or hydrogen Calculate the equivalence ratio for this reaction a. b.
Answer #1


The normal reaction is given as

2H2 + O2 = 2H2O

We need 1 part hydrogen and 6 parts oxygen by mass

mass of hydrogen(2H2) = 4g

mass of oxygen(O2) = 32 g

So, let oxygen available be 4*6=24g

So, 24/32       = 0.75 O2

To make it equivalent,

8H2 + 3O2 = 4H20 +O2

(B) Equivalance ratio of the reaction is given as

X12H2/X602 |X2H2/XO2

0.72/0.27|0.66/0.33

Equivalance ratio is 1.33

That is

0.66/0.33//192 =